OpenStudy (anonymous):

Consider the reaction 2NH3(g) yields N2(g) + 3H2(g) If the rate of delta[H2]/delta t is 0.030 mol L-1 s-1, then delta[NH3]/delta t is? Could someone please explain how to do this problem.

OpenStudy (aaronq):

All you need to do is take the stoichiometric coefficients into account, write a ratio and solve algebraically. \(\dfrac{1}{3}\dfrac{d[H_2]}{dt}=\dfrac{1}{2}\dfrac{d[NH_3]}{dt}\)

OpenStudy (anonymous):

ok cool. I just wanted to make sure I was going in the right direction.

OpenStudy (aaronq):

good stuff